\underline{1}

Template for META08 Full Paper submissions

F. M. Lastname^{1}, F. M. Lastname^{1}, and F. M. Lastname^{2}

1Author Affiliation, Country 2Author Affiliation, Country contact@email.com

Abstract— This file provides a template for writing papers for the conference. The conference proceedings will be published in an electronic format (CD-ROM). The abstract should summarize the approach and the article’s major contributions, emphasizing the importance and significance of the results. It should not exceed 500 words.

1. INTRODUCTION

It is expected that authors will submit carefully written and proofread material. Spelling and grammatical errors, as well as language usage problems, are not acceptable. There is no strict limitation to the number of pages, but it is suggested that the paper length should not exceed 10 pages.

2. SECTION 1

Equations, figures, tables and references should follow a sequential numerical scheme in order to ensure a logical development of subject matter.

For equations,

∇ × \overline{E}(\overline{r}, t) = −\frac{∂\overline{B}(\overline{r}, t)}{∂t}  (1)

Figures and Tables: Give each figure and table a separate concise caption. It begins with Figure or Table, followed by the appropriate Arabic numeral and period. The maximum width of figure or table may not exceed 17 cm. Make symbols, line thickness, and lettering in proper scale in relation to the overall figure size. For best publishing quality, we recommend the EPS format for figures.

3. OTHER SECTIONS ...

Other Sections ...

4. CONCLUSION

Conclusions should state concisely the most important propositions of the paper as well as the author’s views of the practical implications of the results.

ACKNOWLEDGMENT

If any, including all funding information, should be gathered into a brief statement in one paragraph at the end.

REFERENCES

1. Lastname1, F. M., “Title of the journal paper,” Journal Title Abbreviation, Vol. 34, No. 10, 1064–1076, 1986.

2. Lastname1, F. M. and F. M. Lastname2, “Title of the journal paper,” Journal Title Abbreviation, Vol. 34, No. 10, 1064–1076, 1986.

3. Lastname1, F. M., F. M. Lastname2 and F. M. Lastname3, “Title of the journal paper,” Journal Title Abbreviation, Vol. 34, No. 10, 1064–1076, 1986.

4. Lastname, F. M., Book Title, Wiley-Interscience, New York, 1986.

5. Lastname, F. M., “Title of the conference paper,” in Proceedings of International Conference, City, Country, August 2003, 1064-1076.